What is the equation of the circle with center at (2,5) passing through (-5,5) and (-1,1)
1
Expert's answer
2020-10-20T17:53:50-0400

The equation of the circle of radius "R" with center at "(a,b)" is "(x-a)^2+(y-b)^2=R^2". In our case, "(x-2)^2+(y-5)^2=R^2". Since the circle passing through "(-5,5)", we have "(-5-2)^2+(5-5)^2=R^2", and consequently "R=7". On the other hand, since the circle passing through "(-1,1)", we have "(-1-2)^2+(1-5)^2=R^2", and therefore "R=5". This contradiction shows that such a circle does not exist.